OpenStudy (anonymous):

Please help me: Prove this expression do not depend on x "Cos^2(pi/4+x) + cos^2(pi/4-x)"

OpenStudy (aum):

\(\large \cos(A) = \sin(\frac \pi2-A)\) \(\large \cos(\frac \pi4+x) = \sin(\frac \pi2-\frac \pi4-x) = \sin(\frac \pi4-x)\) \(\large \cos^2(\frac \pi4+x) + \cos^2(\frac \pi4-x) = \sin^2(\frac \pi4-x) + \cos^2(\frac \pi4-x) = 1\)

OpenStudy (anonymous):

yup that's right!